OpenStudy (anonymous):

Determine the solution set of (x + 1)^2 = 25.

OpenStudy (anonymous):

{-8, 6} {-8, 4} {-6, 4}

OpenStudy (anonymous):

does the right side seems the sqr of something?

OpenStudy (anonymous):

what ?

Parth (parthkohli):

Square root of both sides. \( \color{Black}{\Rightarrow x + 1 = \pm 5 }\) 1st value of x \( \color{Black}{\Rightarrow x + 1 = 5 }\) \( \color{Black}{\Rightarrow x = 4 }\) 2nd value of x. \( \color{Black}{\Rightarrow x + 1 = -5 }\) \( \color{Black}{\Rightarrow x = -6}\)

OpenStudy (anonymous):

thanks parth